OpenStudy (anonymous):

Identify the asymptotes, removable discontinuities, and intercepts for the graph of each function? (will post in a second)

OpenStudy (anonymous):

1) \[\frac{ x^2-12x+20 }{ 3x }\] 2) \[\frac{ x^2-7x-8 }{ x^3+64 }\] 3) \[\frac{ 2x^3-3x+1 }{ x^3-5x+7 }\] 4) \[\frac{ x^2+4x-5 }{ x^2+8x+15 }\]

OpenStudy (anonymous):

soooo yeahhh im totally lost in answering one so i cant do any lol - any help ???